OpenStudy (rosales0228):

WILL GIVE MEDAL AND FAN!!! Solve the equation the square root of the quantity x plus 3 plus 4 equals 5 for the variable. Show each step of your solution process.

OpenStudy (rosales0228):

how do you do that though? I need to show the steps @TrojanPoem

OpenStudy (rosales0228):

would i just then ad the 3 and four together and then subtract if from the 5???

OpenStudy (anonymous):

Mathway.com it could help you

OpenStudy (rosales0228):

Thanks!

OpenStudy (anonymous):

it help you

OpenStudy (rosales0228):

@ncruz yeah it did thanks so much! @TrojanPoem ya but i suck at them!

OpenStudy (trojanpoem):

Sorry,

OpenStudy (trojanpoem):

The square root of x sqrt(x) + 3 + 4= 5

OpenStudy (rosales0228):

It okay. do you know how to find the degree of polynomials?

OpenStudy (rosales0228):

−10p2q + 7p3q − 5q i have to find the degree of this polynomial

OpenStudy (trojanpoem):

Steps to find the variable: 1) Sum both ( 4 + 3) 2) minus 5- the sum of (4 + 3) 3) square both sides

OpenStudy (trojanpoem):

The third degree

OpenStudy (rosales0228):

okay thanks!

OpenStudy (trojanpoem):

-10p^2q + 7p^3 q - 5 q take q as a common factor q * (-10p^2 + 7 p^3 - 5) arrange according to power 7p^3 - 10p^2 - 5

OpenStudy (trojanpoem):

The biggest power is the degree

OpenStudy (rosales0228):

Okay thanks so much again! on the one before that, do i just leave them squared?

OpenStudy (trojanpoem):

sqrt(x) + 7 = 5 sqrt(x) = - 2 x = 4

OpenStudy (rosales0228):

Okay, but when I go to check it they dont match

OpenStudy (trojanpoem):

What do you mean ?

OpenStudy (rosales0228):

when im solving a radical equation i have to check my work so i have to plug 4 back in for x and see if they both equal out to 4

OpenStudy (trojanpoem):

Sqrt(x) + 3 + 4 =5 sqrt(x) = -2 No answers, sqaure roots can't result in -2

OpenStudy (trojanpoem):

@Michele_Laino

OpenStudy (michele_laino):

maybe the right equation is like this: \[\sqrt {x + 3} + 4 = 5\]

OpenStudy (michele_laino):

as we can see, its solution is \(x=-2\)

OpenStudy (trojanpoem):

You're right, the quantity ( x + 3)

OpenStudy (michele_laino):

the radical exists, if and only if the subsequent condition holds: \[x + 3 \geqslant 0\] so, \(x=-2\) is acceptable

OpenStudy (trojanpoem):

Yeah ( sqrt( -2 + 3 ) + 4= 5

OpenStudy (michele_laino):

how to find that solution? first step: I rewrite that equation as below: \[\sqrt {x + 3} = 5 - 4 = 1\] second step: I take the square of both sides, so I can write this: \[x + 3 = 1\] please simplify